OpenStudy (anonymous):

Find the value of 8 + 2^2 (-3 + 9) ÷ 3.

OpenStudy (anonymous):

First simplify it a bit 1. 2^2 = 4 2. (-3+9) = 6 3. and you get 8 + 4* 6 / 3 4. simplify again 4*6 = 24 5. and you get 8 + 24/3 6. 8 + 24/3 = 8+ 8 7. 8 + 8 = 16 You need to follow your order of operations, multiply , divide, then add and subtract